OpenStudy (anonymous):

Find the range given the domain: y= 6 over x + 1 Domain: {-2, 3, 6}

jimthompson5910 (jim_thompson5910):

When x = -2, y = 6/(x+1) = 6/(-2+1) = 6/(-1) = -6 So when x = -2, y = -6 When x = 3, y = 6/(x+1) = 6/(3+1) = 6/(4) = 3/2 So when x = 3, y = 3/2 When x = 6, y = 6/(x+1) = 6/(6+1) = 6/7 So when x = 6, y = 6/7 So the range is {-6, 3/2, 6/7}